[Brahms-l] Shift report 20040212 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Thu Feb 12 2004 - 07:37:49 EST
------------------------------------------------------------
Date: Thu Feb 12 07:37:49 2004, Shift: 00:00-08:00
Supervisor: Ionut ARSENE, with: Radek Karabowicz
Log book#15 pages 53 - 53

Shift Summary:
------------------------------------------------------------
1. Beam for 5.5 hours (thanks god).

2. We taked the runs between 10281 and 10287.
 We finished the half field, A polarity in FS at 4 degrees.
 We did also the D2=D3=D4=0 field with D1 at B861.
   MRS is still at 45 degrees B2255

3. No detector or magnet problems. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 12 07:38:16 2004

This archive was generated by hypermail 2.1.8 : Thu Feb 12 2004 - 07:38:34 EST